Mother made 15 cookies. There are 10 cookies with raisins and 10 cookies with nuts. How many cookies have both cookies and nuts?
Answer by ikleyn(52781)

.
10 + 10 - 15 = 20 - 15 = 5.

For details, see the lesson Counting elements in sub-sets of a given finite set in this site.